Find an equation of the line that passes through the given point and has the given slope. Express your answer in slope-intercept form.
(8, 0), m = -3

Answer by mananth(16946)

The general equation when we have the slope and a point
(y-y1)=m(x-x1)
m=-3
(x1=8
x2=0
y-0=-3(x-8)
y=-3x+24